OpenStudy (anonymous):

Julian bought 15 pounds of kitty litter for his cat. How many ounces of kitty litter did he buy? A 150 B 270 C 240 D 180

OpenStudy (texaschic101):

1 lb = 16 oz set it up like a proportion 1/16 = 15/x --- (1 lb to 16 oz = 15 lbs to x oz) now cross multiply (1)(x) = (16)(15) x = 240 so 15 lbs = 250 oz

OpenStudy (texaschic101):

oops...my bad...meant to put 15 lbs = 240 oz